# Improving Phishing Email Detection Performance of Small Large Language Models

### v1

@@ 5 Limitation
Our method effectively improves the ability of small LLMs in phishing email detection tasks, achieving performance exceeding models like LLaMA-3.1-70B-Instruct on models around 3B in size. However, limited by time and experimental resources, our research still has the following limitations:
  • Dataset: Due to time and computational resource constraints, we only conducted experiments on the SpamAssassin dataset. Additional experiments on other datasets could be added to further demonstrate the effectiveness of our method.
  • Transferability Study: We did not consider the transferability of our method. Given that LLMs with larger parameter sizes may have advantages in transferability, we may need to design methods to further improve the transferability of small LLMs.
@@ 6 Conclusion
In this study, we explore the feasibility of using small LLMs for phishing email detection. While directly using these small LLMs yields poor detection performance, we designed a process to progressively improve their performance on the phishing detection task. By using prompt engineering to standardize the model output format and encouraging the model to output the reasoning behind its analysis, we first obtain formatted answers with explanations. Subsequently, we significantly improved the phishing email detection capabilities of the small LLMs through explanation augmented fine-tuning, achieving performance that matches or even exceeds mainstream machine learning methods and LLMs with tens of times more parameters. Finally, we further improved the detection performance through model ensemble. In summary, our research indicates the possibility of efficiently performing phishing email detection tasks using small LLMs, and we can obtain high-performance and explainable detection results from these small LLMs.

### v2

@@ 5 Limitation
Our method effectively improves the ability of small LLMs in phishing email detection tasks, achieving performance exceeding models like LLaMA-3.1-70B-Instruct on models around 3B in size. However, due to time and experimental resource constraints, our research still has the following limitations:
  • Dataset: We fine-tuned the three LLMs only on SpamAssassin and CEAS_08 datasets. Additional experiments on other datasets could be added to further validate the effectiveness of our method.
  • Transferability: While we examined the transferability of our method, we did not further propose methods to improve the model’s performance across various datasets.
@@
  • Ensemble Strategy: Beyond confidence ensemble and simple majority voting, more advanced methods, such as weighted voting based on model confidence or validation performance, may further improve performance but were not explored in this work.
@@ 6 Conclusion
In this study, we explore the feasibility of using small LLMs for phishing email detection. While directly using these small LLMs yields poor detection performance, we designed a process to improve their performance on the phishing detection task. By using prompt engineering to standardize the model output format and encouraging the model to output the reasoning behind its analysis, we first obtain formatted answers with explanations. Subsequently, we significantly improved the phishing email detection capabilities of the small LLMs through explanation augmented fine-tuning, achieving performance that matches or even exceeds mainstream machine learning methods and LLMs with tens of times more parameters. Furthermore, we found that it is possible to further improve the detection performance through model ensemble. Finally, our experiments show that fine-tuned small LLMs retain strong performance across multiple unseen datasets, outperforming traditional models and approaching standard-sized LLMs in terms of transferability. In summary, our research indicates the possibility of efficiently performing phishing email detection tasks using small LLMs, and we can obtain high-performance and explainable detection results from these small LLMs.